# GazeHTA: End-to-end Gaze Target Detection with Head-Target Association

### v2

@@ 4.5 Comparison with State-of-the-Art
The quantitative comparison between previous gaze target detection approaches and GazeHTA on the GazeFollow and VideoAttentionTarget datasets is summarized in Tab. 1. It can be seen from the table that GazeHTA outperforms prior works across all evaluation metrics. In contrast to conventional approaches where the understanding of heads and gaze targets is typically acquired independently before integration [9, 13, 2], GazeHTA provides a more unified solution. The advancements stress the importance of concurrently learning the association between heads and gaze targets. Compared to multi-person end-to-end approaches [41, 40], GazeHTA particularly demonstrates significant improvements on mAP, which is specifically designed for measuring head-target association. This shows the effectiveness in accurately identifying both individuals and their associated gaze targets, further highlighting the superiority of GazeHTA in modeling the nuanced correlation between heads and gaze targets.
@@
It is noteworthy that the VideoAttentionTarget dataset is known for its complexity due to the high number of images containing multiple head-target annotations. This characteristic presents a challenge for accurately modeling interactions between multiple individuals and gaze targets. In contrast, the test set of the GazeFollow dataset only contains one in-frame head-target instance per image. Compared to the current SOTA [40], the improvements achieved by GazeHTA on the VideoAttentionTarget dataset (3% in AUC, 32% in Dist., and 26% in mAP) are relatively higher than those on the GazeFollow dataset (1% in AUC, 14% in Avg. Dist., and 12% in mAP). It explicitly demonstrates the efficiency of the integration of the proposed visual association and head feature re-injection techniques.
4.6 Comparison with Diffusion-based Models
Text prompts GazeFollow VideoAttentionTarget AUC Avg. Dist. Min. Dist. mAP AUC Dist. AP mAP VPD [47] Fixed 0.915 0.079 0.038 0.517 0.942 0.082 0.999 0.719 VPD [47] ClipCap [19] 0.920 0.072 0.032 0.556 0.937 0.076 1.000 0.717 VPD [47] - 0.927 0.067 0.028 0.572 0.943 0.083 0.999 0.706 DDP [14] - 0.919 0.077 0.036 0.535 0.932 0.090 0.999 0.685 GazeHTA - 0.935 0.062 0.025 0.639 0.951 0.069 0.999 0.762
One of the contributions of GazeHTA is its utilization of the foundation model Stable Diffusion [27] for the gaze target detection task. To demonstrate its superiority over naive diffusion-based pipelines, we present a comparative analysis between GazeHTA and two SOTA diffusion-based methods, VPD [47] and DDP [14], both of which are based on Stable Diffusion [27] and designed for visual perception tasks, such as semantic segmentation and depth estimation. The results are shown in Tab. 2.
The first baseline, VPD [47], employs a denoising U-Net to extract multi-scale features. We adapt VPD for gaze target detection by taking the scene image as input and then predicting the head heatmaps, gaze heatmaps, and out-of-frame flags. We evaluate the method with three different text prompt options, including i) a fixed text prompt “human faces looking at faces or objects” applied across all input images, ii) text prompts are generated for each image using an image captioner ClipCap [19], and iii) without text prompts and corresponding cross-attention layers.
The second baseline, DDP [14], is a “noise-to-prediction” conditional generative method. We adapt this framework by utilizing a scene image as the condition to iteratively eliminate noise from the input noise, followed by the prediction of head heatmaps, gaze heatmaps, and out-of-frame flags. In our experiment, we set the number of iterations during evaluation to five. The results are shown in the fourth row of Tab. 2.
By comparing the results in Tab. 2 and Tab. 1, it is evident that VPD and DDP achieve comparable or even better results to previous gaze target detection methods on both GazeFollow and VideoAttentionTarget datasets. This shows the potential of pre-trained diffusion models, trained on large-scale datasets, to generalize effectively to the gaze target detection task. Despite involving a multi-step diffusion process, the DDP implementation does not outperform VPD which only uses U-Net as a feature extractor. This observation suggests that the complex diffusion process may not offer benefits for the gaze target detection task. Based on the VPD method with variant text prompt settings, we argue that the benefit of using text prompts for gaze target detection is not significant and also not necessary. Therefore, we remove the text prompt for the final GazeHTA.
Still, GazeHTA outperforms VPD and DDP on both datasets, clearly demonstrating the efficacy of the methodological design presented in this paper for improving performance in the gaze target detection task.
4.7 Ablation Study
Backbone GazeFollow VideoAttentionTarget AUC Avg. Dist. Min. Dist. mAP AUC Dist. AP mAP ✗ ✗ 0.927 0.067 0.028 0.572 0.944 0.081 0.998 0.759 Stable Diffusion [27] ✓ ✗ 0.927 0.068 0.030 0.606 0.946 0.071 1.000 0.751 ✓ ✓ 0.935 0.062 0.025 0.639 0.951 0.069 0.999 0.762 DETR [4] ✓ ✓ 0.928 0.064 0.028 0.607 0.938 0.080 1.000 0.678
4.7.1 Model components.
@@ 4.7 Ablation Study
We further investigate the incremental effectiveness of individual model components within GazeHTA, namely the head feature re-injection and the connection map. The results of this ablation study are presented in the first three rows of Tab. 3, with the third row showing the performance of our complete model GazeHTA. Note that the baseline model without the head feature re-injection nor the connection map is in the first row, which is essentially the VPD [47] without text prompt in Tab. 2.
@@
We consistently observe improvements in both average gaze point distances and mAP over the baseline when incorporating both the connection map and head feature re-injection. On the GazeFollow dataset, we observe improvements of 7% (0.062 vs. 0.067) in Avg. Dist. and 12% (0.639 vs. 0.572) in mAP. On the VideoAttentionTarget dataset, 15% (0.069 vs. 0.081) in Dist. is achieved. However, the incorporation of the connection map alone contributes significantly less to the average gaze point distance on the GazeFollow dataset and mAP on the VideoAttentionTarget dataset. This discrepancy can be attributed to the VideoAttentionTarget dataset containing a larger number of images than GazeFollow with multiple head-gaze instance annotations, which necessitates improvements in head priors. Overall, integrating head feature re-injection and the connection map in GazeHTA yields superior performance across all evaluation metrics compared to its variants.
4.7.2 Backbones.
@@
The results of this experiment are presented in the last row of Tab. 3. In comparison to the state-of-the-art methods presented in Tab. 1, our approach with the DETR backbone consistently outperforms across all evaluation metrics. More importantly, since we use the same backbone as in [41, 40], the improvements in performance solely come from the head feature re-injection and the connection map. This demonstrates the ability of the proposed components to generalize beyond a specific backbone.
Compared to GazeHTA that employs Stable Diffusion as the backbone (third row in Tab. 1), using DETR as the backbone achieves inferior performance. It may be attributed to the fact that DETR is trained with a limited number of object classes, whereas Stable Diffusion is trained to generate realistic images, necessitating the learned features to encompass an understanding of diverse objects present in the training data.
4.8 Qualitative Results
@@
In Fig. 4, we present a visual comparison between Tonini et al. [40] and GazeHTA. We show the detected head bounding boxes alongside their corresponding predicted gaze heatmaps. In the first three columns, Tonini et al. [40] exhibits a limitation in detecting gaze targets when objects are not correctly identified in the object detection stage. In contrast, GazeHTA shows much better results with its fully end-to-end framework and the implicit object understanding inherited from the pre-trained diffusion model. The last two columns are examples of multi-person gaze target detection. Similarly, GazeHTA successfully associates all the head-target instances, while Tonini et al. [40] either fails to identify out-of-frame gaze targets or incorrectly associates head-target instances.
5 Discussion

### v3

@@ IV-E Comparison with State-of-the-Art
The quantitative comparison between previous gaze target detection approaches and GazeHTA on the GazeFollow and VideoAttentionTarget datasets is summarized in Tab. I. It can be seen from the table that GazeHTA outperforms prior works across most of the evaluation metrics with large margins. In contrast to conventional approaches where the understanding of heads and gaze targets is typically acquired independently before integration [25, 34, 26, 36], GazeHTA provides a more unified solution. The advancements stress the importance of concurrently learning the association between heads and gaze targets. Note it is not a strictly fair comparison between our method and these two-stage approaches since these methods use the ground-truth head bounding boxes as input and there is no bipartite matching in the evaluation. Compared to end-to-end approaches [27, 28], GazeHTA particularly demonstrates significant improvements on mAP, which is specifically designed for measuring head-target association. This shows the effectiveness in accurately identifying both individuals and their associated gaze targets, further highlighting the superiority of GazeHTA in modeling the nuanced correlation between heads and gaze targets.
@@
It is noteworthy that the VideoAttentionTarget dataset is known for its complexity due to the high number of images containing multiple head-target annotations. This characteristic presents a challenge for accurately modeling interactions between multiple individuals and gaze targets. In contrast, the test set of the GazeFollow dataset only contains one in-frame head-target instance per image. Compared to the current end-to-end SOTA [28], the improvements achieved by GazeHTA on the VideoAttentionTarget dataset (3% in AUC, 37% in Dist., and 27% in mAP) are relatively higher than those on the GazeFollow dataset (0.75% in AUC, 12% in Avg. Dist., and 9% in mAP). It explicitly demonstrates the efficiency of the integration of the proposed visual association and head feature re-injection techniques.
IV-F Ablation Study
Backbone | GazeFollow | VideoAttentionTarget | ||||||||
AUC | Avg. Dist. | Min. Dist. | mAP | AUC | Dist. | AP | mAP | |||
Stable Diffusion [31] | ✗ | ✗ | 0.910 | 0.084 | 0.042 | 0.500 | 0.951 | 0.079 | 1.000 | 0.759 |
✓ | ✗ | 0.931 | 0.064 | 0.027 | 0.601 | 0.956 | 0.066 | 1.000 | 0.753 | |
✗ | ✓ | 0.928 | 0.068 | 0.029 | 0.567 | 0.937 | 0.092 | 1.000 | 0.769 | |
✓ | ✓ | 0.929 | 0.063 | 0.025 | 0.626 | 0.957 | 0.064 | 1.000 | 0.772 | |
DETR [47] | ✓ | ✓ | 0.928 | 0.064 | 0.028 | 0.607 | 0.938 | 0.080 | 1.000 | 0.678 |
IV-F1 Model components.
@@ IV-F Ablation Study
We further investigate the incremental effectiveness of individual model components within GazeHTA, namely the head feature re-injection and the connection map. The results of this ablation study are presented in the first four rows of Tab. II, with the forth row showing the performance of our complete model GazeHTA. Note that the baseline model without the head feature re-injection nor the connection map is in the first row, which is essentially the VPD [33] without the text prompt.
@@
We consistently observe improvements in both average gaze point distances and mAP over the baseline when incorporating both the connection map and head feature re-injection. On the GazeFollow dataset, we observe improvements of 25% (0.063 vs. 0.084) in Avg. Dist. and 24% (0.626 vs. 0.500) in mAP. On the VideoAttentionTarget dataset, 18% (0.064 vs. 0.079) in Dist. is achieved. In general, we can see that employing either the connection map or the head feature re-injection can boost the performances across metrics except few exceptions. Overall, integrating both head feature re-injection and the connection map in GazeHTA yields superior performance compared to its variants.
IV-F2 Backbones.
@@
The results of this experiment are presented in the last row of Tab. II. In comparison to the state-of-the-art methods presented in Tab. I, our approach with the DETR backbone consistently outperforms across all evaluation metrics. More importantly, since we use the same backbone as in [27, 28], the improvements in performance solely come from the head feature re-injection and the connection map. This demonstrates the ability of the proposed components to generalize beyond a specific backbone.
Compared to GazeHTA that employs Stable Diffusion as the backbone (forth row in Tab. I), using DETR as the backbone achieves inferior performance. It may be attributed to the fact that DETR is trained with a limited number of object classes, whereas Stable Diffusion is trained to generate realistic images, necessitating the learned features to encompass an understanding of diverse objects present in the training data.
IV-G Qualitative Results
@@
V Discussion